H.T. Tan Company is preparing the annual financial statements dated December 31 of the current year. Ending inventory informaAssume that a retailers beginning inventory and purchases of a popular item during January included (1) 470 units at $8.70 iAssume that a retailers beginning inventory and purchases of a popular item during January included (1) 470 units at $8.70 iHamilton Company uses a periodic inventory system. At the end of the annual accounting period, December 31 of the current yeaGrants Corporation prepared the following two income statements (simplified for illustrative purposes): First Quarter $ 12,00Required: 1. What effect did this error have on the combined pretax income of the two quarters? 2. Which quarters or quarterRequired: 1. What effect did this error have on the combined pretax income of the two quarters? 2. Which quarters or quarterRequired information [The following information applies to the questions displayed below.) Givoly Inc. uses a periodic inventRequired information [The following information applies to the questions displayed below.] Givoly Inc. uses a periodic invent
Answer #1

As per Chegg policy only one question should be posted and answered at a time. Please upload other questions seperately. Below answering the very first question:

Item Quantity Total Cost Total NRV Lower of Cost or NRV
A 58 $    1,334 $    1,160 $                 1,160
B 88 $    3,344 $    4,224 $                 3,344
C 18 $    1,008 $    1,080 $                 1,008
D 78 $    2,574 $    2,964 $                 2,574
E 358 $    6,444 $    4,654 $                 4,654
Total Inventory Value $ 14,704 $ 14,082 $              12,740
Workings:
Per unit Total
Item Quantity Cost NRV Cost NRV Lower of Cost or Market
(a) (i) (ii) (iii) = (a)*(i) (iv) = (a)*(ii) Lower of (iii) or (iv)
A 58 $          23 $          20 $                 1,334 $           1,160 $            1,160
B 88 $          38 $          48 $                 3,344 $           4,224 $            3,344
C 18 $          56 $          60 $                 1,008 $           1,080 $            1,008
D 78 $          33 $          38 $                 2,574 $           2,964 $            2,574
E 358 $          18 $          13 $                 6,444 $           4,654 $            4,654
Total Inventory Value $              14,704 $        14,082 $          12,740